OpenStudy (shelby1290):

describe the relation in words (0,0),(2,-10),(4,-20),(5,-25)

OpenStudy (anonymous):

y=x-12,y=x-24, y=x-30

OpenStudy (shelby1290):

how do i put that into words?

OpenStudy (shelby1290):

or will that be my answer...

OpenStudy (shelby1290):

@chash1232014 ^^^

OpenStudy (shelby1290):

are you there?

OpenStudy (anonymous):

yes

OpenStudy (shelby1290):

okay so that whole equation is my answer?

OpenStudy (shelby1290):

right?